# ERX: A Fast Real-Time Anomaly Detection Algorithm for Hyperspectral Line Scanning

### v3

@@ Abstract
Detecting unexpected objects (anomalies) in real time has great potential for monitoring, managing, and protecting the environment. Hyperspectral line-scan cameras are a low-cost solution that enhance confidence in anomaly detection over RGB and multispectral imagery. However, existing line-scan algorithms are too slow when using small computers (e.g. those onboard a drone or small satellite), do not adapt to changing scenery, or lack robustness against geometric distortions. This paper introduces the Exponentially moving RX algorithm (ERX) to address these issues, and compares it with existing RX-based anomaly detection methods for hyperspectral line scanning. Three large and more complex datasets are also introduced to better assess the practical challenges when using line-scan cameras (two hyperspectral and one multispectral). ERX is evaluated using a Jetson Xavier NX compute module, achieving the best combination of speed and detection performance. This research paves the way for future studies in grouping and locating anomalous objects, adaptive and automatic threshold selection, and real-time field tests. The datasets and the Python code are available at https://github.com/WiseGamgee/HyperAD.
@@ I Introduction
Detecting anomalies in real time offers an advantage for applications that require fast and autonomous decision making. This has great potential in hyperspectral imaging for remote sensing; a rapidly growing area of research used to monitor, manage, and protect the environment [5, 6, 7, 8, 9, 10]. Hyperspectral imaging, also known as imaging spectroscopy, provides greater confidence than RGB or multispectral imaging when detecting anomalies. This is because the pixels in a hyperspectral image (HSI) have more detailed spectral profiles that uniquely represent the material present [11], as shown in Figure 1. The greater depth of spectral information can make it easier to separate anomalies from the background [1].
This study focuses specifically on anomaly detection using hyperspectral line-scan cameras, as they are an increasingly popular choice of camera for remote sensing applications. Line-scan cameras capture one line of pixels at a time and require the motion of an attached platform such as a drone, aircraft, or satellite to collect an image (Figure 2). They have relatively high signal-to-noise ratios [12], high spatial and spectral resolutions [13], and are easier and cheaper to build with commercially available components [14, 15]. Real-time anomaly detection applications that use these cameras include search and rescue [16], defence surveillance and reconnaissance [17], and hazard detection in agricultural harvesting [18].
@@
However, none of the aforementioned algorithms are designed for real-time processing, let alone for the additional complexities of line scanning. Real-time line scanning is a causal time series process, accessing only the current line of pixels captured by the camera at time $t$ and previous lines [1, 49]. Most HAD methods analyse an image of the entire scene at once, and are only demonstrated on powerful CPUs and GPUs. This requirement for high computational and memory resources makes them unsuitable for applications onboard platforms with limited computing power. In general, real-time HAD is a less explored field that has an established need for more research [19, 50, 51, 52].
Variants of the RX algorithm have emerged as a popular choice for line-scan anomaly detection due to their effectiveness and lightweight design. They use different techniques to process the hyperspectral data pixel-by-pixel or line-by-line, making them ideal for real-time applications. One technique is the Woodbury matrix identity, which is used to recursively update the inverse correlation matrix Mahalanobis distance calculations [49, 53, 54]. Calculating the inverse is an otherwise slow and memory-intensive operation if performed for each new line. Other works apply Cholesky decomposition and forward substitution to avoid calculating the inverse covariance altogether [55, 56].
Window-based methods were used to improve anomaly detection by considering only the local surroundings of each pixel [57, 58, 59, 60, 61, 62], and kernel methods were found to enhance detection by better mapping the relationships between hyperspectral bands [63]. Algorithms using dimensionality reduction have also been proposed, using real-time versions of principal component analysis (PCA) and subspace projection [18, 64]. These reduce the number of spectral bands to improve the processing speed. However, even with these techniques, line scan anomaly detection algorithms lack a combination of three key characteristics:
@@
    HyperAD: https://github.com/WiseGamgee/HyperAD, an openly available Python repository that contains the implementation of ERX and other algorithms from the literature. This is the first open-source repository for real-time anomaly detection for hyperspectral line scanning, which promotes accessibility and future research.
@@ II-D Local/Window Algorithms
Window-based algorithms extract anomalies locally rather than at a global level, improving detection performance [57] and adaptability. As shown in Figure 3, a standard window is a small region surrounding the current pixel being processed. Double-windows discard a small inner region of pixels, minimising the contribution of the same object to the local background statistics.
@@ II-G Summary
Existing methods employ techniques such as the Woodbury matrix identity, Cholesky decomposition, local windows, kernel functions, and PCA to enhance the RX algorithm for line scanning. However, they miss three crucial aspects: scalability for processing large numbers of HSI bands and pixels in real time, adaptability to changing scenery, and robustness to geospatial distortions. Previous work focused on small, corrected image datasets with limited scenery variation, revealing additional gaps in evaluating line-scanning algorithms under practical conditions. In the following section, we present ERX and three larger and more diverse datasets to address these challenges. Two datasets have changing scenery to test adaptability, with one also having geometric distortions to better evaluate robustness.
@@ IV-D Experiments
All of the experiments are conducted on an NVIDIA Jetson Xavier NX device, a compact edge computer ideal for applications onboard small platforms such as a drone. This device features 8GB of RAM, a 6-core Carmel ARMv8.2 CPU, and is selected to operate in the 20W 6-core power mode. Real-time anomaly detection is simulated in a Python environment using a sampler and a detector. The sampler sequentially feeds each line (from left to right in the HSI datasets) to the detector (the algorithm being tested).
@@ IV-D3 Ablation Study
This study analyses how altering or removing sparse random projection and the exponentially moving mean and covariance affect ERX’s performance. The first part adjusts the number of dimensions ($d$) reduced by sparse random projection: from 1 to 50 for the beach and synthetic datasets, and from 1 to 10 for the SNP dataset (since it only has 13 bands). Sparse random projection is also completely removed for all datasets. ROC curves will be generated for all datasets, and the speed will also be assessed because of its dependence on the number of dimensions. Only AUCs and speeds are provided for the synthetic dataset, as it presents the greatest detection challenge and changing $d$ has a consistent impact on speed regardless of the dataset.
The second part of the ablation study investigates the ERX momentum parameter $\alpha$, which modifies ERX’s detection performance based on the size of the anomaly. Different values are tested between $\alpha=1$ (using only the mean and covariance of the current line for detection) and $\alpha=1e-4$ (heavily weighing past data). This component is also eliminated by using incremental mean and covariance (IMC) updates. The implementation of the line-by-line IMC follows a batch version of Welford’s algorithm for online mean and covariance updates [81], which weighs all previous lines equally.
@@ V-A Speed Comparison
Figure 9: A comparison of the algorithms’ average speed on the Jetson device across five iterations. The top graph shows speed versus the number of bands, and the bottom graph shows speed versus the number of pixels per line. The y-axis (speed) is on a log-10 scale, and the shaded areas represent the standard deviation.
@@ V-C1 Sparse Random Projection
TABLE III: Ablation Study - Mean AUC $\pm$ SD with Varying Momentum ($\alpha$).
@@
V-C2 Momentum
The detection results for varying momentum across all datasets are given in Table III, and Table IV compares the average speeds when using exponential moving averages (EMAs) and when they are removed. Anomaly heatmaps of the synthetic dataset in Figure 13 visually demonstrate how momentum affects ERX’s detection performance.
TABLE IV: Ablation Study - Speed (Lines per Second) With and Without Momentum ($\alpha$).
Model Beach Synthetic SNP
No EMAs $527$ $535$ $574$
EMAs $\bm{573}$ $\bm{548}$ $\bm{599}$
(a) $\alpha=1$
(b) $\alpha=0.1$
(c) $\alpha=1e-3$
(d) No EMAs
Adjusting the momentum of ERX demonstrates that it improves the detection performance in various scenarios (Table III). ERX works well with high momentum on the beach and synthetic datasets, but performance declines after $\alpha<0.01$. This suggests that when $\alpha\leq 1e-3$, ERX is slow to adapt and over-emphasises past pixels. Small $\alpha$ values overly bias the model toward initial background estimates given by the mean and covariance of the first line, further slowing the adaptation to changing scenery. The synthetic dataset and the flipped version are most affected, as the high anomaly rate further skews the background mean and covariance estimates. Figure 13 illustrates this, with anomalies blending into the background after the first two columns of anomalous targets. ERX performs consistently across different momentum values on the SNP dataset, as expected due to more uniform scenery and infrequent anomalies. Momentum values of 0.1 and 0.01 generally perform best across all datasets, but it is clear that adjusting the momentum can enhance anomaly detection based on the size and frequency of the anomalies in the scene.
@@ V-D Limitations & Future Work
Creating an open-source Python repository improves accessibility and reproducibility for future projects, but at the cost of slower processing speeds. Implementing these algorithms in faster languages like C++ would significantly improve their processing speeds. If faster languages were used, more algorithm types, such as those using kernel and window methods, may become feasible for real-time applications. Performing field tests for extended periods, beyond simulations, is crucial to validating the practicality of the ERX algorithm.
@@ VI Conclusion
ERX performed best overall, yet the datasets were challenging and open up some promising pathways for future work. Some opportunities include field tests to validate algorithm practicality, adaptive and automatic threshold selection, and unsupervised and self-supervised algorithms that group anomalous pixels into objects and locate them.

### v4

@@ Abstract
Detecting unexpected objects (anomalies) in real time has great potential for monitoring, managing, and protecting the environment. Hyperspectral line-scan cameras are a low-cost solution that enhance confidence in anomaly detection over RGB and multispectral imagery. However, existing line-scan algorithms are too slow when using small computers (e.g. those onboard a drone or small satellite), do not adapt to changing scenery, or lack robustness against geometric distortions. This paper introduces the Exponentially moving RX algorithm (ERX) to address these issues, and compares it with four existing RX-based anomaly detection methods for hyperspectral line scanning. Three large and more complex datasets are also introduced to better assess the practical challenges when using line-scan cameras (two hyperspectral and one multispectral). ERX is evaluated using a Jetson Xavier NX edge computing module (6-core CPU, 8GB RAM, 20W power draw), achieving the best combination of speed and detection performance. ERX was 9 times faster than the next-best algorithm on the dataset with the highest number of bands (108 band), with an average speed of 561 lines per second on the Jetson. It achieved a 29.3% AUC improvement over the next-best algorithm on the most challenging dataset, while showing greater adaptability through consistently high AUC scores regardless of the camera’s starting location. ERX performed robustly across all datasets, achieving an AUC of 0.941 on a drone-collected hyperspectral line scan dataset without geometric corrections (a 16.9% improvement over existing algorithms). This work enables future research on the detection of anomalous objects in real time, adaptive and automatic threshold selection, and real-time field tests. The datasets and the Python code are openly available at https://github.com/WiseGamgee/HyperAD, promoting accessibility and future work.
@@ I Introduction
Detecting anomalies in real time is essential for applications requiring rapid insights and fast decision making. This enables immediate anomaly mapping when surveying a region of interest, and onboard system responses to detected anomalies. Onboard detection can also reduce data transmission requirements by sending only key insights rather than raw, multi-channel image data. These capabilities are particularly valuable in hyperspectral imaging for remote sensing, a rapidly growing area of research used to monitor, manage, and protect the environment [5, 6, 7, 8, 9, 10]. Hyperspectral imaging, also known as imaging spectroscopy, provides greater confidence than RGB or multispectral imaging when detecting anomalies. This is because the pixels in a hyperspectral image (HSI) have more detailed spectral profiles that uniquely represent the material present [11], as shown in Figure 1. The greater depth of spectral information can make it easier to separate anomalies from the background [1].
This study focuses specifically on anomaly detection using hyperspectral line-scan cameras, as they are an increasingly popular choice of camera for remote sensing applications. Line-scan cameras capture one line of pixels at a time and require the motion of an attached platform such as a drone, aircraft, or satellite to collect an image (Figure 2). They have relatively high signal-to-noise ratios [12], high spatial and spectral resolutions [13], and are easier and cheaper to build with commercially available components [14, 15]. Real-time line scanning is a causal time-series operation that processes each line of pixels as they are received at time $t$, with access only to previous lines [1, 16]. Some real-time anomaly detection applications that use these cameras include search and rescue, defence surveillance and reconnaissance, and hazard detection in agricultural harvesting [17, 18, 19].
@@
However, none of the aforementioned algorithms are designed for real-time processing, let alone for the additional complexities of line scanning. Most HAD methods analyse the entire image at once, and are only demonstrated on powerful CPUs and GPUs. This requirement for high computational and memory resources makes them unsuitable for real-time applications onboard platforms with limited computing power. In general, real-time HAD is a less explored field that has an established need for more research [20, 50, 51, 52].
Variants of the RX algorithm have emerged as a popular choice for line-scan anomaly detection due to their effectiveness and lightweight design. They use different techniques to process the hyperspectral data pixel-by-pixel or line-by-line, making them ideal for real-time applications. One technique is the Woodbury matrix identity, which is used to recursively update the inverse correlation matrix for the Mahalanobis distance calculations [16, 53, 54]. Calculating the inverse is an otherwise slow and memory-intensive operation if performed for each new line. Other works apply Cholesky decomposition and forward substitution to avoid calculating the inverse covariance altogether [55, 56].
Window-based methods were used to improve anomaly detection by considering only the local surroundings of each pixel [57, 58, 59, 60, 61, 62, 63], and kernel methods were found to enhance detection by better mapping the relationships between hyperspectral bands [64]. Algorithms using dimensionality reduction have also been proposed, using real-time versions of principal component analysis (PCA) and subspace projection [19, 65]. These reduce the number of spectral bands to improve the processing speed.
However, even with these techniques, line scan anomaly detection algorithms lack a combination of three key characteristics:
@@
    HyperAD: https://github.com/WiseGamgee/HyperAD, an openly available Python repository that contains the implementation of ERX and other algorithms from the literature. This is the first open source repository for real-time anomaly detection for hyperspectral line scanning, which promotes accessibility and future work.
@@ II-D Local/Window Algorithms
Window-based algorithms extract anomalies locally rather than at a global level, improving detection performance [57, 58] and adaptability. As shown in Figure 3, a standard window is a small region surrounding the current pixel being processed. Double-windows discard a small inner region of pixels, minimising the contribution of the same object to the local background statistics.
@@ II-G Summary
Existing methods employ techniques such as the Woodbury matrix identity, Cholesky decomposition, local windows, kernel functions, and dimensionality reduction to enhance the RX algorithm for line scanning. However, they were found to lack one or more key requirements: scalability for processing large numbers of HSI bands and pixels in real time, adaptability to changing scenery, and robustness to geospatial distortions. Previous work also focused on small, corrected image datasets with limited scenery variation, revealing an additional gap in evaluating line-scanning algorithms under practical conditions. In the following section, we present ERX and three larger and more diverse datasets to address these challenges. Two datasets have changing scenery to test adaptability, with one also having geometric distortions to better evaluate robustness.
@@ IV-D Experiments
All of the experiments are conducted on an NVIDIA Jetson Xavier NX device, a compact edge computer ideal for real-time applications onboard small platforms such as a drone. This device features 8GB of RAM, a 6-core Carmel ARMv8.2 CPU, and is selected to operate in the 20W 6-core power mode. Real-time anomaly detection is simulated in a Python environment using a sampler and a detector. The sampler sequentially feeds each line from left to right in the HSI datasets, to the detector (the algorithm being tested).
@@ IV-D3 Ablation Study
This study analyses how altering or removing sparse random projection and the exponentially moving mean and covariance affect ERX’s performance. The first part adjusts the number of dimensions ($d$) reduced by sparse random projection: from 1 to 50 for the beach and synthetic datasets, and from 1 to 10 for the SNP dataset (since it only has 13 bands). ROC curves will be generated for all datasets, and the speed will also be assessed because of its dependence on the number of dimensions. Only AUCs and speeds are provided for the synthetic dataset, as it presents the greatest detection challenge and changing $d$ has a consistent impact on speed regardless of the dataset.
The second part of the ablation study investigates the ERX momentum parameter $\alpha$, which modifies ERX’s adaptability, and detection performance based anomaly size and frequency. Different values are tested between $\alpha=1$ (using only the mean and covariance of the current line for detection) and $\alpha=1e-4$ (heavily weighing past data). This component is also eliminated by using incremental mean and covariance (IMC) updates. The implementation of the line-by-line IMC follows a batch version of Welford’s algorithm for online mean and covariance updates [82], which weighs all previous lines equally.
@@ V-A Speed Comparison
Figure 9: A comparison of the algorithms’ average speed on the Jetson device across five iterations. The top graph shows speed versus the number of bands, and the bottom graph shows speed versus the number of pixels per line. The y-axis (speed) is on a log-10 scale, and the shaded areas represent the standard deviation. Note: the drastic speed change in the top graph occurs when the number of bands exceeds 95, due to the OpenBLAS library switching threading algorithms.
@@ V-C1 Sparse Random Projection
V-C2 Momentum
The detection results for varying momentum across all datasets are given in Table IV, and Table III compares the average speeds when using exponential moving averages (EMAs) and when they are removed. Anomaly heatmaps of the synthetic dataset in Figure 13 visually demonstrate how momentum affects ERX’s detection performance.
TABLE III: Ablation Study - Speed (Lines per Second) With and Without Momentum ($\alpha$).
Model Beach Synthetic SNP
No EMAs $527$ $535$ $574$
EMAs $\bm{573}$ $\bm{548}$ $\bm{599}$
(a) $\alpha=1$
(b) $\alpha=0.1$
(c) $\alpha=1e-3$
(d) No EMAs
TABLE IV: Ablation Study - Mean AUC $\pm$ SD with Varying Momentum ($\alpha$).
@@ V-C2 Momentum
Adjusting the momentum of ERX demonstrates that it improves the detection performance in various scenarios (Table IV). ERX works well with high momentum on the beach and synthetic datasets, but performance declines after $\alpha<0.01$. This suggests that when $\alpha\leq 1e-3$, ERX is slow to adapt and over-emphasises past pixels. Small $\alpha$ values overly bias the model toward initial background estimates given by the mean and covariance of the first line, further slowing the adaptation to changing scenery. The synthetic dataset and the flipped version are most affected, as the high anomaly rate further skews the background mean and covariance estimates. Figure 13 illustrates this, with anomalies blending into the background after the first two columns of anomalous targets. ERX performs consistently across different momentum values on the SNP dataset, as expected due to more uniform scenery and infrequent anomalies. Momentum values of 0.1 and 0.01 generally perform best across all datasets, but it is clear that adjusting the momentum can enhance anomaly detection based on the size and frequency of the anomalies in the scene.
@@ V-D Limitations & Future Work
Creating a Python repository improves accessibility and reproducibility for future projects, but at the cost of potentially slower processing speeds. Implementing these algorithms in faster languages like C++ could improve their speed. With faster processing, more algorithm types, such as those that use kernel and window methods, may become feasible for real-time applications. Performing field tests for extended periods, beyond simulations, is also crucial for validating the practicality of the ERX algorithm.
@@ VI Conclusion
ERX performed best overall, yet the datasets were challenging and open up some promising pathways for future work. These opportunities include field tests to validate algorithm practicality, adaptive and automatic threshold selection, and unsupervised and self-supervised algorithms that group anomalous pixels into objects and locate them.